OpenStudy (anonymous):

indefinite integral (π/2sinnx) dx from 0,π

OpenStudy (anonymous):

either it is indefinite or is from 0 to PI but both cannot be. Can you be a little more specific. Is it\[\frac{ \pi }{ 2 } \int\limits_{0}^{\pi}\sin(nx)dx\] what you are looking for?

OpenStudy (anonymous):

∫π/2.sin(x)dx integration

OpenStudy (anonymous):

Which function gives sin(x) as derivative?

OpenStudy (anonymous):

u know Fourier series in that

OpenStudy (anonymous):

Yes, I do.